OpenStudy (anonymous):

Discuss and identify any possible restrictions that exist with (or in) the resulting rational expression.

OpenStudy (anonymous):

\[\frac{ 2x }{ x+4 }-\frac{ x }{ x-4 }\]

OpenStudy (anonymous):

@abdela25

OpenStudy (anonymous):

@bookworm00981 @Preetha @tester97

OpenStudy (abdela25):

@whpalmer4

OpenStudy (confusionist):

Okay, let's just say that mass tagging is not helpful, so cool it. If you want a reply, wait instead of spamming our notifications.

OpenStudy (anonymous):

@Nateyboo @GeekyChick @nrg100 @orion92898 @KingKhan94 @shamil98 @pinkmommy @SeaTurtle113 sorry i was trying to get people that could help i was not sure if what the answer would be

OpenStudy (anonymous):

oh ok

OpenStudy (justnick09):

ahh i see my notifcation r gonna blow up...yipee

OpenStudy (anonymous):

well we see that if x = 4 or -4 in the denominator, then that causes the answer to be undefined. so x is restricted by 4 and -4. Because you can't divide by 0

OpenStudy (anonymous):

oh ok thanks for explaining

OpenStudy (anonymous):

but do you think you could help me solve it @ilovemathbutneedhelp

OpenStudy (anonymous):

So first we have to find a common denominator.

OpenStudy (anonymous):

ok and that would be 4 right

OpenStudy (anonymous):

the common denominator for the problem would be (x+4)(x-4) so we are able to combine the numerators

OpenStudy (anonymous):

so it would be \[\frac{ 2 }{ (x+4)(x-4)}\]

OpenStudy (anonymous):

okay this is a tricky problem, so i will show you step by step here: http://www.twiddla.com/1503747

OpenStudy (radar):

X can not = 4, or a -4 do you see why?

OpenStudy (radar):

When addition is completed you will have a resulting fraction where the degree of the numerator and the degree of the denominator are equal (both second degree) and since both the coefficients are the asymptote is a horizontal line y=1

OpenStudy (anonymous):

yes i see why now i have another problem and i ghad tagged you in it so could you help me there

OpenStudy (radar):

Lets see, I am not good in many areas of math. show me.

OpenStudy (radar):

Oh yes,that problem was not an equation to be solved, but just an expression that some characteristics they wanted.

OpenStudy (radar):

Where is the problem

OpenStudy (anonymous):

i had tagged you i will do it again